OpenStudy (anonymous):

Kahlil said that using the equation (a + b)^2 = a^2 + 2ab +b^2, he can find a similar equation for (a - b)^2. Examine his argument. If it is correct, supply any missing steps, or justifications; if it is incorrect point out why.

Parth (parthkohli):

\[(a + b)^2 = a^2 + 2ab + b^2 \]\[(a - b)^2 = (a + (-b))^2 = a^2 + 2a(-b) + (-b)^2 \]

OpenStudy (anonymous):

Can I get an explaination please.

OpenStudy (anonymous):

Replace +b with -b and see what happens.

OpenStudy (anonymous):

Substraction is defined as adding the negative of a number. So, we simply replace the minus sign by adding the negative of b.